Friendlys Quick Loans, Inc., offers you five for seven or I knock on your door. This means you get $5 today and repay $7 w
Answer #1

Amount borrowed = $5
Amount repaid = $7
Period = 1 week

Weekly interest rate = (Amount repaid / Amount borrowed) - 1
Weekly interest rate = ($7 / $5) - 1
Weekly interest rate = 1.40 - 1
Weekly interest rate = 0.40

Answer a.

Annual interest rate = 52 * Weekly interest rate
Annual interest rate = 52 * 0.40
Annual interest rate = 20.80 or 2080.00%

Answer b.

Effective annual rate = (1 + Weekly interest rate)^52 - 1
Effective annual rate = (1 + 0.40)^52 - 1
Effective annual rate = 39,687,875.8299 - 1
Effective annual rate = 39,687,874.8299 or 3,968,787,482.99%